[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon May 01, 2017

Jeffrey Lewitsky - NOAA Federal jeffrey.lewitsky at noaa.gov
Mon May 1 20:48:45 UTC 2017


Chris,

How do we run this script?

The 18z run for the Atlantic/NH2 just finished but only the first 24
hours came into GFE.

Jeff

On 5/1/17, Christopher Mello - NOAA Federal <christopher.mello at noaa.gov> wrote:
> All,
> *nwps update*. I wrote a script that will re-ingest  the nwps data a second
> time into GFE after the nwps finishes running. This will help ensure all
> the available  hours gets ingested into GFE...
>
> If the hourly data is still missing...then it is likely an upstream issue
> with the model out side of AWIPS2. We will observe the nwps data in GFE
> this week to find out.
>
> Chris
>
> On Mon, May 1, 2017 at 9:52 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
>
>> Here is today's summary of NHC computer operations:
>> --- Mello
>>
>> ----------
>>
>> *TAFB:*
>>
>>
>> TAFB:
>> 1. TAFB reported no 00z Atlantic/NH2 NWPS data ingested into D2D/GFE
>> beyond 60 hours. This problem continues to randomly occur. Also, no 21z
>> SREF SBN data ingested into D2D.
>>
>> Follow up. The 00Z and 06Z nwps model runs ingested fully overnight
>> Sunday
>> for both domains. The problem might be EDEX getting overwhelmed at
>> certain
>> times (EDEX congestion might also cause the reported SREF ingest
>> failure).
>>
>> Possible Solution. Mello will create an nwps Desktop icon today. Clicking
>> on this icon will reingest all the current AWIPS2 nwps model data into
>> GFE.
>> This script will determine if the problem is an EDEX ingest issue. If the
>> data does not update beyond 60 hours, then the nwps model run is failing
>> before it gets into AWIPS2.
>>
>> 2. Matt Sardi continues to work with NCO to resolve the NAWIPS
>> performance
>> issues. Matt identified some X configuration errors on both the
>> TILEFISH'S
>> and BALLYHOO's computers. Forecasters reported that TILEFISH was stable
>> but
>> sluggish at times over the weekend.
>> *HSU:*
>>
>> HSU:
>> 1. TSB/NCO will swap out the last old ATCF monitor in HSU this week(The
>> Pacific Desk).
>>
>> 2. NCO, Joe Zajic (HQ) and Monica Bozeman are still troubleshooting
>> multiple issues with GOES-16 data flow. Expect sporadic data outages
>> while
>> developers clean up the software.
>> *Other items:*
>>
>> *NHC RFCs:*
>>
>>
>>
>
>
> --
> Chris Mello
> National Hurricane Center
> TSB
>


More information about the Ncep.nhc.nco_contacts mailing list